OpenStudy (anonymous):

Which equation represents the line passing through the point (4, −16) and is perpendicular to the line y equals negative two thirds x plus 8? y equals two thirds x minus 22 y equals negative two thirds x minus 22 y equals negative three halves x minus 22 y equals three halves x minus 22

OpenStudy (anonymous):

\[y=-\frac{2}{3}x+8\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok first question is do you know what the slope of \(y=-\frac{2}{3}x+8\) is ?

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

ok good next question is do you know what the slope of a perpendicular line will be?

OpenStudy (anonymous):

if not, say "idk" and i will tell you

OpenStudy (anonymous):

no idea, sorry

OpenStudy (anonymous):

ok fine

OpenStudy (anonymous):

it is the "negative reciprocal" in other words flip it and change the sign

OpenStudy (anonymous):

so since the line has slope \(-\frac{2}{3}\) the perpendicular line will have slope \(\frac{3}{2}\)

OpenStudy (anonymous):

so 3/2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that should eliminate some answers right away, but we can go ahead and find the equation for the line using the point - slope formula do you know that one?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok it is \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

wait, y=mx+b?

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

no, that is the "slope intercept" form

OpenStudy (anonymous):

put \[y-y_1=m(x-x_1)\] make a direct substitution with \[x_1=4,y_1=-16,m=\frac{3}{2}\]

OpenStudy (anonymous):

you get \[y-(16)=\frac{3}{2}(x-4)\] or \[y+16=\frac{3}{2}(x-4)\]

OpenStudy (anonymous):

then multiply out on the left to get \[y+16=\frac{3}{2}x-6\] so \[y=\frac{3}{2}x-22\]

OpenStudy (anonymous):

i hope that is one of the choices!

OpenStudy (anonymous):

it is thank you so much!